OpenStudy (anonymous):

Solve the equation using factoring x2 + 7x + 12 = 0 A. {2, 6} B. {-2, -6} C. {3, 4} D. {-3, -4}

OpenStudy (anonymous):

alright, what two number add up to 7 and multiplies to 12

OpenStudy (anonymous):

?

OpenStudy (anonymous):

none

OpenStudy (anonymous):

what's 4+3?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

what's 4*3?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

so the answer is c ?

OpenStudy (anonymous):

(x+4)(x+3)=0

OpenStudy (anonymous):

x+4=0 x=-4 x+3=0 x=-3